Dilution Formula. If you take 2.0 liters of soluti
Answer #1

To answer this we need the Molarity of solution from part a. However, lets assume it is 0.1M

Formula to use:

M_{1}V_{1} = M_{2}V_{2}

M1 = 0.1 M

V1 = 2.0 lit

V2 = 2.0 + 1.5 = 3.5 lit

M2 = ?

M_{2} = M_{1}V_{1}/V_{2}

= 0.1 * 2/3.5 = 0.057M